# Paper/main_online_appendix

### v3

@@ Abstract
This paper develops a theory of how tradeoffs govern comparison complexity, and how this complexity generates systematic mistakes in choice. In our model, options are easier to compare when they involve less pronounced tradeoffs, in particular when they are 1) more similar feature-by-feature and 2) closer to dominance. These two postulates yield tractable measures of comparison complexity in the domains of multiattribute, lottery, and intertemporal choice. We show how our model organizes a range of behavioral regularities in choice and valuation, such as context effects, preference reversals, and apparent probability weighting and hyperbolic discounting. We test our model experimentally by varying the strength and nature of tradeoffs. First, we show that our complexity measures predict choice errors, choice inconsistency, and cognitive uncertainty in binary choice data across all three domains. Second, we document that manipulations of comparison complexity can reverse classic behavioral regularities, in line with the predictions of the theory. We apply our theory to study strategic obfuscation by firms in a pricing game.
@@ 1 Introduction
We often face difficult tradeoffs when comparing choice options, and standard economic models abstract from this difficulty. In choosing utility services, households must trade off fixed fees and variable usage fees. In shopping for mortgages, homebuyers weigh brokerage fees and closing costs against interest rates. In selecting health insurance plans, individuals balance premium and deductible costs as well as other financial and non-financial features of coverage. In each of these settings, choosing optimally requires the decision-maker to engage in a cognitively challenging process in the presence of tradeoffs: they must both assess the relative values of different option features, and aggregate value across features.
Tradeoffs are a near-ubiquitous feature of economic decisions. Yet despite evidence from psychology that individuals struggle with navigating tradeoffs (e.g. Tversky,, 1969), economists lack a formal understanding of how tradeoffs make options difficult to compare, and how this difficulty distorts choice. Existing theories of choice complexity, which have focused on modeling what makes options hard to value, or how menu size affects complexity, do not capture the difficulty of tradeoffs. For instance, options like mortgage contracts may be difficult to value, yet are easy to compare in the absence of tradeoffs — i.e. if one transparently dominates the other — and tradeoffs complicate choice even in binary menus. Moreover, existing theories tend to be domain-specific, which limits researchers’ ability to organize a growing body of evidence suggesting that various empirical regularities across different domains of behavioral economics are driven by complexity.
This paper makes four contributions. First, we develop a theory of comparison complexity in the domains of multiattribute, lottery, and intertemporal choice, which formalizes the common principle that comparisons are difficult when they involve pronounced tradeoffs across option features. Second, we show how our theory can rationalize an array of documented biases and instabilities from different choice domains, including context effects, preference reversals, and apparent probability weighting and hyperbolic discounting. Third, we use rich experimental data in all three domains to test the behavioral implications of our theory, including novel predictions on how these choice patterns can be eliminated or even reversed by manipulating the nature of tradeoffs. Finally, we apply our model to draw out implications for strategic obfuscation by firms in product markets.
Theory of comparison complexity. We model a decision-maker (DM) who is uncertain about the values of two options, $x$ and $y$, and chooses in a Bayesian fashion based on a noisy signal on how these values compare. The precision of this signal, $\tau_{xy}$, captures the ease of comparison between $x$ and $y$, and governs the decision-maker’s likelihood of choosing the higher-valued option. We develop a theory of how $\tau$ depends on the features of choice options in multiattribute, lottery, and intertemporal choice.
@@
$\begin{array}[]{ll}x:&\text{\$27 w.p. 25\%, \$3 w.p. 75\%}\\ x^{\prime}:&\text{\$9 for sure}\\ y:&\text{\$20 w.p. 20\%, \$3.2 w.p. 80\%}\\ \end{array}$
@@
$\begin{array}[]{ll}x:&\text{\$60 in 61 days}\\ x^{\prime}:&\text{\$100 in 3 years}\\ y:&\text{\$40 in 60 days}\end{array}$
@@
Our theory is motivated by the idea that decision-makers struggle to aggregate tradeoffs across option features, and that not all comparisons require the same degree of aggregation. Tradeoffs complicate comparisons because decision-makers may be uncertain over the relative importance of utility-relevant features, and even absent preference uncertainty, may find it difficult to aggregate advantages and disadvantages across features. To illustrate, consider the three choice environments in Figure 1. Across these domains, the comparison between $x$ and $y$ is simple — there is little need to make tradeoffs to see that $x$ is better than $y$. On the other hand, the comparison between $x^{\prime}$ and $y$ is less obvious, as the DM must now engage with non-trivial tradeoffs: 1a) involves a tradeoff between the monthly fee vs. usage fee, 1b) involves trading off a higher maximum payout against a lower payout probability, and 1c) involves a tradeoff between payout amounts and delays.
@@
where the numerator contains the value difference between the two options and the denominator is a distance metric measuring their dissimilarity. Intuitively, similar options require less aggregation of tradeoffs to compare, as the DM can divert attention from features that are similar across options and so more easily assess differences. This intuition echoes work in psychology and economics (Tversky and Russo,, 1969; Rubinstein,, 1988), and we follow recent work in stochastic choice (He and Natenzon, 2023a, ) in our formalism.
@@
Multiattribute $U(x)=\sum_{k}\beta_{k}x_{k}$ $\tau_{xy}=H\left(\dfrac{|U(x)-U(y)|}{d_{L1}(x,y)}\right)$ $d_{L1}(x,y)=\displaystyle\sum_{k}|\beta_{k}(x_{k}-y_{k})|$
[…]
@@
This model generates two key implications. First, in binary choice, comparison complexity leads to noisy, but unbiased, choice. Second, in larger menus, comparison complexity generates systematic distortions: a) context effects in multinomial choice, which occur when competing alternatives are hard to compare to each other but differ in their comparability to other options in the menu, and b) “pull-to-center” effects in the valuation of choice options – that is, when the option being valued is hard to compare to prices, valuations are compressed towards the center of the price list.
@@
To build intuition, consider the options $x\succ z^{1}\succ z^{2}\succ...\succ z^{n}$, where $\{z^{1},z^{2}...,z^{n}\}$ are themselves perfectly comparable, but hard to compare to $x$. Figure 2 illustrates the choice environments of interest. First, consider the binary menu in panel a). As the DM receives only an imprecise signal on how $x$ and $z^{1}$ compare, she may err in her choice, but is more likely to choose the better option $x$. Next, consider the addition of $z^{2}$ to the menu in panel b). The DM is uncertain how $x$ compares to either $z^{1}$ or $z^{2}$, but is certain that $z^{1}$ is superior to $z^{2}$: this additional information inflates the DM’s assessment of $z^{1}$ relative to $x$, and so distorts choice toward $z^{1}$, producing a context effect. Note that this information also worsens the DM’s assessment of $z^{2}$ relative to $x$, and so can pull her valuation of $x$ between that of the unambiguously ranked options $z^{1}$ and $z^{2}$. This logic generalizes to valuation tasks as in panel c), where the DM values $x$ against $\{z^{1},z^{2},..,z^{n}\}$ in a multiple price list: the difficulty of comparing $x$ to the unambiguously ranked items in the price list pulls her valuation of $x$ toward the center of that ranking.
(a) Binary Choice
$\begin{array}[]{ll}{x}\,\,\,\,\,\,\,\text{vs.}&\,\,\,{z^{1}}\\ \end{array}$
(b) Context Effects
[…]
(c) Valuations
[…]
Figure 2: Choice environments. $z^{1},z^{2},...,z^{n}$ are easy to compare to each other, but hard to compare to $x$.
@@
These forces rationalize a range of documented behavioral regularities in choice and valuation. In multiattribute choice, the context effects predicted by our model straightforwardly generate familiar decoy/asymmetric dominance effects. Given our lottery and intertemporal complexity measures, which predict which risky and intertemporal prospects are hard to compare to prices, our model also rationalizes documented regularities in valuation: preference reversals and apparent probability weighting and hyperbolic discounting, specifically because these patterns reflect pull-to-center biases.
@@
To illustrate, consider a canonical paradigm used to estimate the probability weighting function: eliciting the certainty equivalents of a simple lottery $l$ that pays \$$\overline{w}$ with probability $p\in[0,1]$. For $p$ sufficiently close to 0 or 1, $l$ has a near-dominance relationship with the certain payments in the price list, so valuations are close to accurate. For interior values of $p$, however, tradeoffs between payoff amounts and probabilities make $l$ harder to compare to the price list, producing pull-to-center distortions: small probabilities are overvalued and large probabilities are undervalued, generating inverse-S probability weighting. The same logic generates apparent hyperbolic discounting in the valuation of delayed payments.
@@
Importantly, our model predicts that these distortions are not generic, but instead arise from the difficulty of comparing options to prices. As such, our model generates the novel prediction that one can reverse these distortions by manipulating tradeoffs in the valuation task. Suppose that instead of valuing lotteries against certain payments, our decision-maker is tasked with valuing certain payments against lotteries: assessing the payoff probability $p$ that makes $l$ indifferent to \$$\$w\in[0,\overline{w}]$ for sure. The pull-to-center distortions that produce classic probability weighting in certainty equivalents now generate the opposite pattern in this paradigm: low certain payments are overvalued and high certain payments are undervalued, causing the agent to appear risk-averse over low probabilities and risk-seeking over high probabilities. Moreover, our model predicts that hyperbolic discounting can also be reversed by similarly inverting the role of the price list and the option being valued.
@@
Since these pull-to-center distortions affect valuations but not direct choice, our model predicts systematic inconsistencies between these two modes of eliciting preferences. As such, our model straightforwardly rationalizes documented preference reversals in lottery choice: cases where subjects favor lower-risk options when choosing between lotteries, yet exhibit the opposite preference in their valuations of those lotteries; through the same logic, the model generates similar reversals in intertemporal choice. The model also predicts that these reversals can be eliminated by manipulating the relative ease of comparing each option to prices – specifically by changing the numeraire against which options are valued.
@@
Experimental tests. We conduct two sets of experiments. First, we test the validity of our proposed complexity measures using three large-scale binary choice datasets corresponding to our domains of interest: choice between induced-values multiattribute goods, lotteries, and time-dated payoff streams. In line with the theory, our complexity measures are strongly predictive of choice inconsistency, choice errors, and subjective uncertainty over choices in all three datasets. When a choice problem involves a lower value-dissimilarity ratio, subjects are more likely to make inconsistent choices across repeat instances of that problem, and are more likely to make choice “errors", where we define an error as a) choosing the lower-value option in multiattribute choice and b) choosing the less-preferred option according to a best-fit model of risk and time preferences in lottery and intertemporal choice, respectively. These relationships are quantitatively meaningful: multiattribute choice inconsistency and error rates are more than four times as large for choice problems with the lowest versus highest value-dissimilarity ratio, with similarly pronounced relationships in our other domains. Furthermore, the value dissimilarity ratio is strongly predictive of the subjective difficulty of comparisons, as measured by the uncertainty subjects report over their choices.
Second, we run a series of experiments to demonstrate that both preference reversals and classic distortions in lottery and intertemporal valuation are in part outgrowths of comparison complexity — specifically by showing that these patterns can be eliminated or reversed by manipulating the nature of tradeoffs as our model predicts.
In our preference reversal experiments, we 1) reproduce classic reversals in lottery choice and document similar reversals in intertemporal choice, and 2) test the novel prediction that these reversals can be eliminated by manipulating the relative ease of comparing options to prices. Consistent with our model and existing findings, we find that subjects give higher valuations for high-risk vs. low-risk lotteries (high-delay vs. low-delay payments) when valuing options against money, yet exhibit the opposite preference when choosing directly between options. In line with the model, however, the inconsistency between valuations and direct choice disappears when subjects instead value lotteries (delayed payments) using probability equivalents (time equivalents).
Finally, we document that two classic valuation patterns, probability-weighting and hyperbolic discounting, can be reversed by manipulating tradeoffs in the valuation task. In line with model predictions, we find that using probability equivalents rather than certainty equivalents to estimate the probability weighting function results in apparent underweighting of small probabilities and overweighting of large probabilities. Similarly, relative to the discounting revealed by their present value equivalents, subjects’ time equivalents exhibit overvaluation for delays close to the present and undervaluation for longer delays.
@@
Benchmarking predictive power. Using our binary choice datasets, we structurally estimate our model and compare its performance to leading benchmarks. In all three domains, our model explains a large share of variation in choice rates not captured by existing models, and substantially outperforms leading models in lottery and intertemporal choice, explaining 10–22% more variation than cumulative prospect theory and hyperbolic discounting. Following Fudenberg et al., (2022), we characterize the completeness of our theory in capturing the predictable variation in choice rates relative to a flexible predictive model trained on our data. Our model explains 70% of the predictable variation in our multiattribute choice data, and over 90% in lottery and intertemporal choice. We also estimate the restrictiveness of our models, a measure of how well the model fits simulated “synthetic” choice data (Fudenberg et al.,, 2023). Our model is less restrictive than the leading model in multiattribute choice, but is more restrictive than leading models in lottery and intertemporal choice.
Application to strategic obfuscation. We use our model to study a Bertrand pricing game in which firms sell products of identical value to consumers, and can influence how comparable their products are to their competitors through choice of how value is distributed across product attributes. We find that in a symmetric environment, firms design dissimilar products to minimize the ease of comparison to their competitors and therefore soften price competition, which leads to spurious differentiation and positive equilibrium markups. On the other hand, when there are vertical cost differences between firms, firms with a sufficiently large cost advantage have incentives to design products that are easy to compare to high-cost competitors, and compete directly on price. As a result, our model predicts that an increase in competitive pressure brought on by lower costs can actually increase prices and obfuscation in equilibrium. We connect our predictions to field evidence, and discuss the extent to which our model of complexity-driven mistakes can be distinguished from alternative accounts of product proliferation, such as preference heterogeneity.
@@
This paper adds to a theoretical literature on complexity by formalizing a distinct source of choice difficulty and establishing its empirical relevance. Existing theories tend to formalize complexity at the option level, modeling what makes options difficult to value (Puri,, 2023; Hu,, 2023; Salant and Spenkuch,, 2022; Gabaix and Laibson,, 2017), or instead study how menu size affects complexity (de Lara and Dean,, 2024; Gerasimou,, 2018). These accounts are incomplete. For instance, two lotteries can be complex in the sense of having many states, and yet easy to compare if one dominates the other; and even in binary menus, choice can be difficult in the presence of tradeoffs. By formalizing how tradeoffs make options difficult to compare, our theory provides a complementary account that can capture these features of complexity.
@@
Our paper also contributes to a growing literature studying the cognitive foundations of economic decision-making, which aims to explain multiple empirical regularities from different choice domains using a set of common cognitive principles. We show how a single mechanism – the difficulty of making tradeoffs – rationalizes phenomena as distinct as probability weighting, hyperbolic discounting over money, preference reversals, context effects, and the large variation of noise in binary choice. Our paper thus relates to work on the cognitive foundations of prospect theory-like behavior (Khaw et al.,, 2021; Frydman and Jin,, 2021; Enke and Graeber,, 2023; Oprea,, 2022) and hyperbolic discounting (Gabaix and Laibson,, 2017; Enke et al.,, 2023). In contrast to the majority of the literature, we develop a model and experimental tests that tie together regularities across different choice domains.
@@
Our theory not only rationalizes these behavioral phenomena, but also makes progress in organizing a body of evidence documenting their instability and context-dependence. Whereas canonical distortions such as probability weighting and hyperbolic discounting are typically modeled as a fixed part of the agent’s value function, experiments have shown that they are unstable across choice contexts: for instance, probability weighting is far less pronounced in direct choice between lotteries and certain payments relative to valuation tasks (Harbaugh et al.,, 2010; Bouchouicha et al.,, 2023). Relatedly, a long-standing literature has documented inconsistencies between preference elicitation methods (see Seidl,, 2002, for a review). These findings pose basic issues for economists hoping to apply insights from behavioral research to a given setting, or to draw inferences from choice data. We propose a theory that organizes these findings, which we validate by testing novel predictions on how manipulating tradeoffs can eliminate or even reverse these choice patterns.
@@
Our model builds on a recent stochastic choice literature (Natenzon,, 2019; He and Natenzon, 2023a, ; He and Natenzon, 2023b, ): in particular, it belongs to a general moderate utility class axiomatized in He and Natenzon, 2023a , wherein binary choice probabilities are a function of the value difference between two options normalized by a distance metric. We make two contributions to this literature. First, we bring this class of models closer to data by proposing specific distance metrics computable in three choice domains. Our measures are characterized by a common set of properties reflecting tradeoff complexity, and in particular respect dominance, unlike existing models in the moderate utility class. Second, we bring experimental evidence to bear on this literature by showing that our theory rationalizes a range of documented choice patterns.
@@
Section 2 develops the measure of comparison complexity. Section 3 develops the multinomial choice model and studies its behavioral implications. Section 4 describes the experimental tests of the model. Section 5 discusses our application to strategic firm obfuscation. Section 6 discusses relationships between our theory and existing models, and Section 7 concludes. Appendix C contains proofs of all results stated in the main text.

### v4

# Cassidy Shubatt    Jeffrey Yang

@@ Abstract
This paper develops a theory of how tradeoffs govern the difficulty of comparisons. In our model, options are easier to compare when they involve less pronounced tradeoffs – when they are 1) more similar feature-by-feature and 2) closer to dominance. These postulates yield tractable measures of comparison complexity in three domains: multiattribute, lottery, and intertemporal choice. Our model rationalizes multiple behavioral regularities, such as context effects, preference reversals, and apparent probability weighting and hyperbolic discounting. In choice data spanning all three domains, our model predicts errors, inconsistency, and cognitive uncertainty. Manipulating tradeoffs reverses classic behavioral regularities, in line with model predictions.
@@ 1 Introduction
We often face difficult tradeoffs when comparing choice options, and standard economic models abstract from this difficulty. In choosing utility services, households trade off fixed fees and variable usage fees. In shopping for mortgages, homebuyers weigh brokerage fees and closing costs against interest rates. In selecting health insurance plans, individuals balance premium and deductible costs as well as other financial and non-financial features of coverage. In each of these settings, choosing optimally requires the decision-maker to engage in a cognitively challenging process in the presence of tradeoffs: they must both assess the relative values of different option features, and aggregate value across features.
Tradeoffs are a feature of many economic decisions. Yet despite evidence from psychology that individuals struggle to navigate tradeoffs (e.g. Tversky,, 1969; Tversky and Shafir,, 1992; Shugan,, 1980), economists lack a formal understanding of how tradeoffs make options hard to compare, and how this difficulty distorts choice. At the same time, a growing body of evidence suggests that well-known behavioral regularities from different choice domains may be partially driven by complexity. Taken together, this raises the question of whether a formal model of tradeoff difficulty can make sense of existing findings and generate new implications for economic behavior.
This paper makes three contributions. First, we develop a theory of comparison complexity in the domains of multiattribute, lottery, and intertemporal choice, which formalizes the principle that comparisons are difficult when they involve pronounced tradeoffs across option features. Second, we show how our theory can rationalize an array of documented biases and instabilities ranging across multiple choice domains, including context effects, preference reversals, and apparent probability weighting and hyperbolic discounting. Third, we use rich experimental data in all three domains to test the behavioral implications of our theory, including novel predictions on how these choice patterns can be eliminated or even reversed by manipulating the nature of tradeoffs.
Theory of comparison complexity. We model a decision-maker (DM) who is uncertain about the values of two options, $x$ and $y$, and chooses based on a noisy signal on how these values compare. The precision of this signal, $\tau_{xy}$, captures the ease of comparison between $x$ and $y$, and governs the decision-maker’s likelihood of choosing the higher-valued option. We develop a theory of how the ease of comparison $\tau_{xy}$ depends on the features of choice options in multiattribute, lottery, and intertemporal choice.
@@
[…]
@@
[…]
@@
Our theory is motivated by the idea that decision-makers struggle to aggregate tradeoffs across option features, and that not all comparisons require the same degree of aggregation. Tradeoffs complicate comparisons because decision-makers may be uncertain over the relative importance of utility-relevant features, and even absent preference uncertainty, may find it difficult to aggregate advantages and disadvantages across features. To illustrate, consider the three choice environments in Figure 1. Across these domains, the comparison between $x$ and $y$ is simple — there is little need to make tradeoffs to see that $x$ is better than $y$. On the other hand, the comparison between $x^{\prime}$ and $y$ is less obvious, as the DM must now engage with non-trivial tradeoffs: 1a) involves a tradeoff between the phone plan monthly fee vs. usage fee, 1b) involves trading off a higher maximum payout against a lower payout probability, and 1c) involves a tradeoff between payout amounts and delays.
@@
where the numerator contains the value difference between the two options and the denominator is a distance metric measuring their dissimilarity. Intuitively, similar options require less aggregation of tradeoffs to compare, as the DM can divert attention from features that are similar across options and so more easily assess differences. This intuition echoes work in psychology and economics (Tversky and Russo,, 1969; Rubinstein,, 1988), and we follow recent work in stochastic choice (He and Natenzon,, 2024) in our formalism.
@@
Multiattribute
$U(x)=\sum_{k}\beta_{k}x_{k}$
$\tau_{xy}=H\left(\dfrac{|U(x)-U(y)|}{d_{L1}(x,y)}\right)$ $d_{L1}(x,y)=\displaystyle\sum_{k}|\beta_{k}(x_{k}-y_{k})|$
Lottery
$EU(x)=\sum_{w}u(w)f_{x}(w)$
$\tau_{xy}=H\left(\dfrac{|EU(x)-EU(y)|}{d_{CDF}(x,y)}\right)$
$d_{CDF}(x,y)=\displaystyle\int_{0}^{1}|u(F^{-1}_{x}(q))-u(F^{-1}_{y}(q))|\,dq$
Intertemporal
$PV(x)=\sum_{t}\delta^{t}x_{t}$
$\tau_{xy}=H\left(\dfrac{|PV(x)-PV(y)|}{d_{CPF}(x,y)}\right)$
$d_{CPF}(x,y)=\ln\left(\frac{1}{\delta}\right)\displaystyle\int_{0}^{\infty}\delta^{t}|M_{x}(t)-M_{y}(t)|\,dt$
@@
This model generates two key implications. First, in binary choice, comparison complexity leads to noisy, but unbiased, choice. Second, in larger menus, comparison complexity generates systematic distortions: a) context effects in multinomial choice, which occur when competing alternatives are hard to compare to each other but differ in their comparability to other options in the menu, and b) “pull-to-center” effects in the valuation of choice options – that is, when the option being valued is hard to compare to prices in a multiple price list, valuations are compressed towards the center of the price list.
@@
To build intuition, consider the options $x\succ z^{1}\succ z^{2}$, where $\{z^{1},z^{2}\}$ are themselves perfectly comparable, but hard to compare to $x$. First, consider the binary choice $x$ vs. $z^{1}$. As the DM receives only an imprecise signal on how $x$ and $z^{1}$ compare, she may err in her choice, but is more likely to choose the better option $x$. Next, consider the addition of $z^{2}$ to the menu: $x$ vs. $z^{1}$ vs $z^{2}$. The DM is uncertain how $x$ compares to either $z^{1}$ or $z^{2}$, but is certain that $z^{1}$ is superior to $z^{2}$: this additional information inflates the DM’s assessment of $z^{1}$ relative to $x$, and so distorts choice toward $z^{1}$, producing a context effect. Note that this information also worsens the DM’s assessment of $z^{2}$ relative to $x$, and so can pull her valuation of $x$ between that of the unambiguously ranked options $z^{1}$ and $z^{2}$. This logic generalizes to valuation tasks, where the DM values $x$ against $\{z^{1},z^{2},..,z^{n}\}$ in a multiple price list: the difficulty of comparing $x$ to the unambiguously ranked items in the price list pulls her valuation of $x$ toward the center of that ranking.
@@
These forces rationalize a range of documented behavioral regularities. In multiattribute choice, the context effects predicted by our model straightforwardly generate familiar decoy/asymmetric dominance effects. Given our lottery and intertemporal complexity measures, which predict which risky and intertemporal prospects are hard to compare to prices, our model can rationalize documented regularities in valuation: preference reversals and apparent probability weighting and hyperbolic discounting, specifically because these patterns can be generated or exaggerated by a pull-to-center bias in valuation.
@@
To illustrate, consider a canonical paradigm used to estimate risk preferences: eliciting the certainty equivalents of a simple lottery $l$ that pays $\overline{w}$ with probability $p\in[0,1]$. For $p$ sufficiently close to 0 or 1, $l$ has a near-dominance relationship with the certain payments in the price list, so valuations are close to accurate. For interior values of $p$, however, tradeoffs between payoff amounts and probabilities make $l$ harder to compare to the price list, producing pull-to-center distortions: small probabilities are overvalued and large probabilities are undervalued, generating inverse-S probability weighting. The same logic can generate apparent hyperbolic discounting in the valuation of delayed payments.
@@
Importantly, our model predicts that these patterns are not generic, but instead arise from the difficulty of comparing options to prices. As such, it predicts that one can reverse these distortions by manipulating tradeoffs in the valuation task. Suppose that instead of valuing lotteries against certain payments, our decision-maker is tasked with valuing certain payments against lotteries: assessing the payoff probability $p$ that makes $l$ indifferent to $w\in[0,\overline{w}]$ for sure. The pull-to-center distortions that produce classic probability weighting in certainty equivalents now generate the opposite pattern in this paradigm: low certain payments are overvalued and high payments are undervalued, causing the agent to appear relatively risk-averse (risk-seeking) over low (high) probabilities, consistent with experimental work comparing the two valuation methods (Sprenger,, 2015; Feldman and Ferraro,, 2024). Our model also makes the novel prediction that hyperbolic discounting can be similarly reversed by inverting the role of the price list and the option being valued.
@@
Since these pull-to-center distortions affect valuations but not direct choice, our model predicts apparent inconsistencies between these preference elicitation modes. Our model straightforwardly rationalizes documented preference reversals in lottery choice: cases where subjects favor lower-risk options when choosing between lotteries, yet exhibit the opposite preference in their lottery valuations. The model also predicts that these reversals can be eliminated by manipulating the relative ease of comparing each option to prices – specifically by changing the numeraire against which lotteries are valued, consistent with findings from experimental manipulations (Butler and Loomes,, 2007; Slovic et al.,, 2002). Again, the model makes analogous novel predictions for intertemporal preference elicitations.
@@
Experimental tests. We conduct two sets of experiments. First, we test the validity of our proposed complexity measures using three large-scale binary choice datasets in our domains of interest: choice between induced-values multiattribute goods, lotteries, and time-dated payoff streams. In line with the theory, our complexity measures are strongly predictive of choice inconsistency, choice errors, and subjective uncertainty over choices in all three datasets. When a choice problem involves a lower value-dissimilarity ratio, subjects are more likely to make inconsistent choices across repeat instances of that problem, and are more likely to make choice “errors", where we define an error as a) choosing the lower-value option in multiattribute choice and b) choosing the less-preferred option according to a best-fit preference model in lottery and intertemporal choice. These relationships are quantitatively meaningful: multiattribute choice inconsistency and error rates are more than four times as large for choice problems with the lowest versus highest value-dissimilarity ratio, with similarly pronounced relationships in our other domains. Furthermore, the value-dissimilarity ratio is strongly predictive of subjects’ reported uncertainty over their choices.
@@
In all three domains, our model explains a large share of variation in choice rates not captured by existing models, and substantially outperforms leading models in lottery and intertemporal choice by 10-22%. We also conduct completeness and restrictiveness exercises (Fudenberg et al.,, 2022, 2023). In terms of completeness, our model explains 70% of the predictable variation in multiattribute choice, and over 90% in lottery and intertemporal choice. In terms of restrictiveness, our model is less restrictive than the leading model in multiattribute choice, but more restrictive than leading models in lottery and intertemporal choice, suggesting that our gains in predictive power do not come at substantial costs to parsimony relative to benchmark models.
Second, we run a series of experiments to demonstrate that both preference reversals and classic distortions in lottery and intertemporal valuation are in part driven by comparison complexity — specifically by showing that these patterns can be eliminated or reversed by manipulating tradeoffs as our model predicts. In our preference reversal experiments, we 1) reproduce classic reversals in lottery choice and document similar reversals in intertemporal choice, and 2) show that these reversals can be eliminated by manipulating the relative ease of comparing options to prices. In line with the model, the documented inconsistency between valuations and direct choice disappears when subjects instead value lotteries (delayed payments) using probability equivalents (time equivalents).
Finally, we document that two classic valuation patterns, probability-weighting and hyperbolic discounting, reverse when we manipulate tradeoffs in the valuation task. In line with model predictions and existing findings, we find that using probability equivalents rather than certainty equivalents to estimate risk preferences results in apparent underweighting of small probabilities and overweighting of large probabilities. Similarly, relative to the discounting revealed by their present value equivalents, subjects’ time equivalents exhibit overvaluation for delays close to the present and undervaluation for longer delays.
@@
This paper adds to a theoretical literature on complexity by formalizing a distinct source of choice difficulty and establishing its empirical relevance. Existing theories tend to emphasize what makes individual options difficult to value (Puri,, 2025; Hu,, 2023; Salant and Spenkuch,, 2022; Gabaix and Laibson,, 2017), or study how menu size affects complexity (de Lara and Dean,, 2024; Gerasimou,, 2018). These accounts do not speak to the difficulty of assessing tradeoffs, which plays an important role in choice complexity. For instance, two lotteries can be complex in the sense of having many states, and yet easy to compare if one dominates the other; and even in binary menus, choice can be difficult in the presence of tradeoffs. By formalizing a notion of tradeoff difficulty, our theory provides a complementary account that can capture these features of complexity.
@@
Our paper also contributes to a growing literature studying the cognitive foundations of economic decision-making, which aims to explain multiple empirical regularities from different choice domains using a set of common cognitive principles. We show how a single mechanism – the difficulty of making tradeoffs – can rationalize phenomena as distinct as probability weighting, hyperbolic discounting over money, preference reversals, context effects, and the large variation of noise in binary choice. Our paper thus relates to work on the cognitive foundations of prospect theory-like behavior (Khaw et al.,, 2021; Frydman and Jin,, 2021; Enke and Graeber,, 2023; Oprea,, 2024) and hyperbolic discounting (Gabaix and Laibson,, 2017; Enke et al.,, 2025). In contrast to the majority of the literature, we develop a model and experimental tests that tie together regularities across different choice domains.
@@
Our theory not only rationalizes these behavioral phenomena, but also makes progress in organizing a body of evidence documenting their instability and context-dependence. Whereas canonical distortions such as probability weighting and hyperbolic discounting are typically modeled as a fixed part of the agent’s value function, experiments have shown that they are unstable across choice contexts: for instance, probability weighting is far less pronounced in direct choice between lotteries and certain payments relative to valuation tasks (Harbaugh et al.,, 2010; Bouchouicha et al.,, 2023). Relatedly, a long-standing literature has documented inconsistencies between preference elicitation methods (see Seidl,, 2002, for a review). These findings pose basic issues for economists hoping to apply insights from behavioral research to a given setting, or to draw inferences from choice data. We propose a theory that organizes these findings, which we validate by testing model predictions on how manipulating tradeoffs can eliminate or even reverse these choice patterns.
@@
Our model builds on a recent stochastic choice literature (Natenzon,, 2019; He and Natenzon,, 2024, 2023) studying the moderate utility class – models wherein binary choice probabilities are a function of the value difference between two options normalized by a distance metric – as well as research on more general forms of heteroskedastic noise in choice under risk (e.g. Hey,, 1995; Buschena and Zilberman,, 2008; Loomes,, 2005). We make two contributions to these literatures. First, we propose specific moderate utility representations in three choice domains, and show how these representations are characterized by a common set of properties reflecting tradeoff complexity. Second, we formalize how our form of noise can rationalize systematic biases in richer choice settings beyond binary choice, and bring experimental evidence to bear on these predictions.
@@
Section 2 develops the measure of comparison complexity. Section 3 develops the multinomial choice model and studies its implications. Section 4 describes the experimental tests of the model. Section 5 discusses relationships between our theory and existing models, and Section 6 concludes. Appendix C contains proofs of all results stated in the main text.